OpenStudy (anonymous):

Solve: –2x – 13 = 8x + 7

OpenStudy (lgbasallote):

im thinking you have a specific question for this...?

OpenStudy (anonymous):

ADD 2x to both sides AT first

OpenStudy (anonymous):

then subtract 7

OpenStudy (anonymous):

-13 =10x +7

OpenStudy (anonymous):

from both sides

OpenStudy (anonymous):

Now subtract 7 from both sides

OpenStudy (anonymous):

as @kiamonstaa said

OpenStudy (anonymous):

-2x-8x=7+13 -10x= 20 x= 20/-10 x=-2

OpenStudy (anonymous):

-13-7 divided by 10x?

OpenStudy (anonymous):

well -13-7 would be -20

OpenStudy (anonymous):

I get it!

OpenStudy (anonymous):

thanks guys :D

OpenStudy (anonymous):

welcome